A simple TEX example

This is a very simple example of using TEX.

This text will be in italic. And this is slanted. There are many other fonts available, for instance the typewriter mode which some people regars as a good way of separating computer printouts from normal text.

National characters such as ñ or é can be produced relatively easily. Finnish text containing ä and ö can be also be produced using TEX, and there are special tools for doing it comfortably. Unfortunately the systems do not fully support word division in Finnish; some corrective work by the user is required.

Now to mathematics. You can insert mathematical expressions like (a + b)^{2} or x_{1} + … + x_{n} into the text.

You can also make mathematical expressions appear as separate paragraphs. This is normal for large expressions and equations like

e^{x} = ∑_{i=0}^{∞} \frac{x^{i}}{i!}

Let’s end our small example with some special symbols that can be produced in the mathematical mode: α♡⋆ ↦ ∥. Nice, isn’t it?